Garlic is very common in our kitchen and taste when we eat food. It makes the food delicious and give an aroma that excite you to eat. Allicin is the compound responsible for its smell. The molecular formula of a allicin is C6 H10 OS2


a. What is the molar mass of allicin?

b. How many moles of allicin are present in 8.00 mg of this substance?

c. How many molecules of allicin are in 4.00 mg of this substance

d. How many S atoms are present in 2.00 mg of allicin?


1
Expert's answer
2022-02-09T07:29:02-0500

a.   molar mass = sum of molar masses of individual atoms present.

                   = (6 C × 12.01 + 10 H × 1.008 + 1 O × 15.999 + 2 S × 32.066) g/mol

                    = 162.271 g/mol

Therefore, molar mass of allicin is 162.271 g/mol

b.   Number of moles of allicin = mass of allicin in grams/ molar mass of allicin

               First convert 8.00 mg allicin into grams (g).

          8.00 mg allicin = 8.00 mg allicin × 1g allicin/1000 mg allicin

                                   = 8.00 × 10-3 g allicin

Number of moles of allicin = 8.00 × 10-3 g/162.271 g/mol

                                          = 4.93 × 10-5 moles.

c.   4.00 mg allicin = 4.00 mg allicin × 1g allicin/1000 mg allicin

                       = 4.00 × 10-3 g

 Moles of allicin = 4.00 × 10-3 g/162.271 g/mol = 2.47 × 10-5 moles

2.47 × 10-5 moles × 6.022 × 1023 molecules of allicin/1 mole allicin

                            = 1.49 × 1019 molecules of allicin

d.   2.00 mg allicin = 2.00 mg allicin × 1g allicin/1000 mg allicin

                       = 2.00 × 10-3 g allicin

Mass of S in 2.00 × 10-3 g allicin = [(2×32.066)/162.271] × 2.00 × 10-3 g

                                          = 7.9 × 10-4 g of S

 moles of S = mass of S (g) / molar mass of S

                   = 7.9 × 10-4 g/32.066 g/mol

                    = 2.46 × 10-5 moles of S

 

2.46 × 10-5 moles S × 6.022 × 1023 atoms of S /1 mole S

                            = 1.48 × 1019 atoms of S